OpenStudy (anonymous):

If x ml pure alcohol is added to 400 ml of a 15% solution and after adding its strength becomes 32%, then the value of x is (A) 50 ml (B) 75 ml (C) 200 ml (D) 100 ml

OpenStudy (jackellyn):

400 ml of a 15% solution has 60 ml pure alchohal and 340 ml of water. when x ml of pure alchohal is added, strength becomes 32%, hence (60+x)/(400+x) = 32/100= 8/25 1500+25x = 3200+8x 17x = 1700 x= 100 ml